OpenStudy (anonymous):

Find the average rate of change of the function f over the given interval. f(x) = 0.25x^4 - x^2 - 8x + 3 from x = - 2 to x = 8 I got the answer of 886/10 which is not right.

OpenStudy (anonymous):

it will be \[\frac{ f \left( 8 \right)-f \left( -2 \right) }{ 8-\left( -2 \right) }\]